Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Mike and Emily need to build 2 identical houses. Mike . . .

Expert replies
by Vincen » Tue Jan 09, 2018 6:09 am
Mike and Emily need to build 2 identical houses. Mike, working alone, can build a house in 6 weeks. Emily, working alone, can build a house in 8 weeks. To determine who will do the building they will roll a fair six-sided die. If they roll a 1 or 2, Mike will work alone. If they roll a 3 or 4, Emily will work alone. If they roll a 5 or 6, they will work together and independently. What is the probability both houses will be completed after 7 weeks?

A) 0
B) 1/3
C) 1/2
D) 2/3
E) 1

B is the OA.

I am confused here. I don't know how to start. With the probability or the rates. Experts, can you show me how to solve this PS question. I'd really like to see the answer.
Join the discussion
Source: — Problem Solving |

by DavidG@VeritasPrep » Tue Jan 09, 2018 9:31 am
Vincen wrote:Amy and Emily need to build 2 identical houses. Amy, working alone, can build a house in 6 weeks. Emily, working alone, can build a house in 8 weeks. To determine who will do the building they will roll a fair six-sided die. If they roll a 1 or 2, Amy will work alone. If they roll a 3 or 4, Emily will work alone. If they roll a 5 or 6, they will work together and independently. What is the probability both houses will be completed after 7 weeks?

A) 0
B) 1/3
C) 1/2
D) 2/3
E) 1

B is the OA.

I am confused here. I don't know how to start. With the probability or the rates. Experts, can you show me how to solve this PS question. I'd really like to see the answer.
Executive decision: I'm changing the names here.

First, rephrase the question.

If Amy can do 1 house in 6 weeks, she can do 2 in 12 weeks. If Emily can do 1 house in 8 weeks, she can do 2 in 16 weeks. Neither alone can do 2 houses in 7 weeks.

Together: Amy's rate is 1/6 and Emily's rate is 1/8. Their combined rate is 1/6 + 1/8 = 14/48 = 7/24.
Rate and Time have a reciprocal relationship, so the time to do one house together would be 24/7. Thus, the time to do two houses together would be 48/7. This is less than 7.

So our rephrased question: what is the probability that Amy and Emily work together?

Desired outcomes (work together): Roll a 5 or 6 --> so 2 desired outcomes
Total possible outcomes: 6

Desired/Total = 2/6 = 1/3. The answer is B
Veritas Prep | GMAT Instructor

Veritas Prep Reviews
Save $100 off any live Veritas Prep GMAT Course
Join the discussion